OpenStudy (anonymous):

@inkyvoyd @mathmale @phi PLEASE HELP!!!!!!! A scientist is studying the growth of a particular species of plant. He writes the following equation to show the height of the plant f(n), in cm, after n days: f(n) = 8(1.05)n Part A: When the scientist concluded his study, the height of the plant was approximately 11.26 cm. What is a reasonable domain to plot the growth function? Part B: What does the y-intercept of the graph of the function f(n) represent? Part C: What is the average rate of change of the function f(n) from n = 2 to n = 6, and what does it represent?

OpenStudy (anonymous):

@inkyvoyd @mathmale @phi This is all I have so far and I really need help: :'( Part A: The plant is 11.26 cm tall at the end. How many days? We use 8(1.05)^n=(height of plant)=11.26 To find ‘n’ divide by 8: 8(1.05)^n=11.26 〖1.05〗^n=1.4075

OpenStudy (anonymous):

I don't know why there are question marks, but please ignore them

OpenStudy (phi):

use logs to solve for n \[ 1.05^n = 1.4075 \] take the log of both sides \[ \log(1.05^n) = \log(1.4075) \] now use the useful property \[ \log(a^b) =b \log(a) \]

OpenStudy (anonymous):

How do I find the log?

OpenStudy (phi):

you should get n log(1.05) = log(1.4075) divide both sides by log(1.05) n= log(1.4075)/ log(1.05) it's time to dust off your calculator

OpenStudy (phi):

you can use either log (base 10) or ln (log base 3) it does not matter which. or you can type into google log(1.4075)/ log(1.05)=

OpenStudy (anonymous):

I got 7

OpenStudy (phi):

that means a reasonable domain (of x values) is from 0 days up to 7 days

OpenStudy (phi):

that was part A. can you do part B?

OpenStudy (anonymous):

If the x values are the days, then shouldn't the y values be the height in which it grew for part b?

OpenStudy (phi):

What does the y-intercept of the graph of the function f(n) represent? y-intercept means the y value *on the y-axis* i.e. when x=0 so it's a specific point.

OpenStudy (anonymous):

Part A: Domain will be talking about the values of n. Now when we solve for f(n) = 11.26 we get 7.005... so a reasonable range to plot the graph would be from n = 0 to 8 Part B: The y-intercept is the point where n = 0. But n represents the number of days, i.e. time elapsed. So at n = 0 the point represents the initial reading. Part C: At n = 2, f(2) = 8*(1.05)^2 = 8.82 At n = 6, f(6) = 8*(1.05)^6 = 10.72 The average rate of change is (change in height)/(change in n) The change in height is 10.72 - 8.82 = 1.9 The change in n is 6 - 2 = 4 So the average rate of change is 1.9/4 = 0.475 But what is the average rate of change? It is the average daily (because a single change in n corresponds to a day) growth of the plant (because f(n) represents height) Enjoy :)

OpenStudy (anonymous):

I am slightly lost @phi Apple_pi 's answer seems different from what you said it would be

OpenStudy (phi):

what is confusing ?

OpenStudy (anonymous):

You said that it would be a specific number. apple_pi said that n=0. I don't understand what apple_pi is saying.

OpenStudy (anonymous):

And I definitely don't know if it is the same thing you are saying

OpenStudy (phi):

the y-intercept is the y value when x = 0 or in this case, where they are using n instead of x, where n=0

OpenStudy (anonymous):

The graph will look something like this

OpenStudy (anonymous):

This is the y-intercept

OpenStudy (anonymous):

Okay. So, when n=0, the y intercept is the height in which it has grown in 0 days?

OpenStudy (phi):

they want you to know that the y-value when n=0 is (this is logic or common sense), the height of the plant at day 0

OpenStudy (anonymous):

Yeah, or rather it hasn't grown because it hasn't had time to grow

OpenStudy (anonymous):

I understand now. For C, how do I find the average rate of change? I know it is ƒ(b) − ƒ(a) over b − a

OpenStudy (anonymous):

f(b)-f(a) over b-a

OpenStudy (anonymous):

But how do I figure out what is f(b) and what is f(a) ?

OpenStudy (phi):

use the equation they gave you. apple did the work

OpenStudy (anonymous):

Okay. Thank you!

OpenStudy (phi):

but it's good practice to do it yourself

OpenStudy (anonymous):

That is what I was trying to do

OpenStudy (phi):

you need a calculator (or google)

OpenStudy (anonymous):

I still am having difficulty finding f(a) or f(b). How do I know what is f(A) or f(b)

OpenStudy (phi):

I would do the f(6) - f(2) divided by (6-2) you could also do f(2) - f(6) divided by 2 - 6 (that works, but seems backwards)

OpenStudy (phi):

and f(6) is short-hand for find the definition of f(n), which is f(n) = 8(1.05)^n replace the "n" with a 6, like this f(6) = 8(1.05)^6 and now use a calculator 8 times 1.05 ^ 6

OpenStudy (anonymous):

I got 10.72

OpenStudy (phi):

yes. one down, one to go find f(2)

OpenStudy (anonymous):

8.82

OpenStudy (phi):

yes, now find the difference

OpenStudy (anonymous):

1.9